OpenStudy (briana.img):

A metal worker has several 1-kilogram bars of a metal alloy that contain 23% copper and several 1-kilogram bars that contain 79% copper. How many bars of each type of alloy should be melted and combined to create 48 kilograms of a 44% copper alloy? Set up and solve a system of equations to find the answer.

OpenStudy (briana.img):

What method of system of equations is this?

OpenStudy (briana.img):

@dhirensharma would you mind helping me out?

OpenStudy (anonymous):

yup bring it on!!

OpenStudy (anonymous):

okay soo i will give you the equations in a sec

OpenStudy (anonymous):

btw do you want the solutions or the method that u wanna use ??

OpenStudy (briana.img):

@dhirensharma i want the method and then i'll do it and you'll see if i did it right?

OpenStudy (anonymous):

all right so you can do it by subtitution or elimination method but since this involves decimals you can do it by whatever way u find easier out of these 2 methods

OpenStudy (briana.img):

@dhirensharma oh okay i'll do substitution do you know the equation?

OpenStudy (anonymous):

yup i will have them here fr ya in a sec

OpenStudy (anonymous):

x+y=48 .23x+.79y=12.32

OpenStudy (anonymous):

remember when you are given percents you gotta convert them into decimals by dividing them with 100

OpenStudy (briana.img):

@dhirensharma thank you!! i'll come back to you when i get the answers

OpenStudy (anonymous):

yup i will be waiting

OpenStudy (briana.img):

x+y=48 0.23x+0.79y=12.32 0.23(48-y)+0.79y=12.32 11.04-48y+0.79y=12.32 -48y+0.79y=12.32-11.04 -47.21=1.28 x+-47.21=48 x=95.21

OpenStudy (anonymous):

@briana.img wait sorry i gave u the wrong equation here are the right ones x+y=48 .23x+.79y=21.12

OpenStudy (anonymous):

sorryyyyyyy

OpenStudy (briana.img):

@dhirensharma omg ok but thanks!!

OpenStudy (anonymous):

i am sorry but up i m sure these are absolutely right

OpenStudy (briana.img):

i think i did something wrong but you'll see when i send in the final thing

OpenStudy (anonymous):

okay

OpenStudy (briana.img):

x+y=48 0.23+0.79=21.12 0.23(48-y)+0.79y=21.12 11.04-48y+0.79y=21.12 -48y+0.78y=21.12-11.04 -47.21=10.08

OpenStudy (briana.img):

do i use 10.08 or -47.21 in x+y=48 ?!!?

OpenStudy (anonymous):

okay so i got y=18

OpenStudy (anonymous):

know ur mistake in a sec

OpenStudy (anonymous):

21-6y+5.05y=24.10 11.04-48y+0.79y=21.12 u made the same mistake u made earlier multiplication error

OpenStudy (briana.img):

@dhirensharma where did "21-6y+5.05y=24.10" come from?

OpenStudy (anonymous):

disregard it thats not part of question i sent it accidentally

OpenStudy (anonymous):

so when u were multiplieng by using distributive property u made an error

OpenStudy (anonymous):

do u know your error??

OpenStudy (briana.img):

@dhirensharma not really :(

OpenStudy (anonymous):

okay so i will send you the equations where u made the errror

OpenStudy (briana.img):

ok

OpenStudy (anonymous):

0.23(48-y)+0.79y=21.12 11.04-48y+0.79y=21.12

OpenStudy (anonymous):

let me know if u get it now

OpenStudy (briana.img):

so what was i suppose to do?

OpenStudy (anonymous):

0.23(48-y)+0.79y=21.12 11.04-.23y+.79y=21.12

OpenStudy (briana.img):

@dhirensharma ooooh okay

OpenStudy (anonymous):

0.23y*-y=-0.23y

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

like i said you are pretty good with the concept but be careful while multiplying ;)

OpenStudy (briana.img):

i'm stuck on this -0.23y+0.79y=21.12-11.04

OpenStudy (briana.img):

added -0.23 and 0.79 and got 0.56

OpenStudy (anonymous):

thats right don't frget the y though

OpenStudy (briana.img):

that's right?!?! how?!!?

OpenStudy (briana.img):

thought you said it was 18

OpenStudy (anonymous):

-0.23y+0.79y=.56y

OpenStudy (anonymous):

i meant don't frgt the y at the end of .56

OpenStudy (briana.img):

.56y=10.08?!

OpenStudy (briana.img):

OOOH OKAY

OpenStudy (briana.img):

i learned this method from an example so i guess i didn't fully grasp it

OpenStudy (anonymous):

incase u need more help i can send u a video explaining this stuff

OpenStudy (anonymous):

how about i send it tommorow cause i am in the middle of a summer school so my hands are full today

OpenStudy (briana.img):

@dhirensharma i'm in the middle of summer school too thanks for helping me out!! hopefully i can survive on the next problems!! good luck

OpenStudy (anonymous):

heheeh allright so what school are u in??

OpenStudy (anonymous):

i m in chaparral high

OpenStudy (briana.img):

@dhirensharma oh so you're in vegas?!!?

OpenStudy (anonymous):

nope close to vegas though

OpenStudy (anonymous):

scottsdale AZ

OpenStudy (briana.img):

@dhirensharma i got to ECTA

OpenStudy (briana.img):

@dhirensharma oh there's a school here with the same name

OpenStudy (briana.img):

go*

OpenStudy (anonymous):

thats surprising so are u frm arizona ??

OpenStudy (briana.img):

@dhirensharma no nev ada

OpenStudy (anonymous):

coool place here in AZ it's hot as helllll

OpenStudy (briana.img):

@dhirensharma yeah i just visited arizona i went to the grand canyon lmao

OpenStudy (anonymous):

awesome place isn't it but i bet u got tanned!!?!!?!!

OpenStudy (briana.img):

@dhirensharma just a bit not much :p

OpenStudy (anonymous):

allright so is there any other question u need help with??

OpenStudy (briana.img):

@dhirensharma yeah i just @'d you in this new question